Hello :)
I am working on a worksheet and I really need help with problems such as the following:
Directions:- Use the DISTRIBUTIVE PROPERTY to rewrite the expressions.
13. 9(8 + 7)
14. 11(10 + 5)
16. 16(z + 3)
These are the 3 problems that I really need help with using the Distributive Property.
Please email me back ASAP & Thanks A Lot !! :)
: Hello :)
I am working on a worksheet and I really need help with problems such as the following:
Directions:- Use the DISTRIBUTIVE PROPERTY to rewrite the expressions.
13. 9(8 + 7)
14. 11(10 + 5)
16. 16(z + 3)
These are the 3 problems that I really need help with using the Distributive Property.
Please email me back ASAP & Thanks A Lot !! :)

Answer by stanbon(18787)

Directions:- Use the DISTRIBUTIVE PROPERTY to rewrite the expressions.
You are going to distribute the multiplication over (around) the addition.
13. 9(8 + 7) = 9*8 + 9*7 = 96 + 63 = 159
14. 11(10 + 5) = 11*10 + 11*5 = 110 + 55 = 165
16. 16(z + 3) = 16z + 16*3 = 16z + 48
===============================
Cheers,
Stan H.